# LLM-Powered Agents for Navigating Venice’s Historical Cadastre

### v1

@@ 5.1 Method
5.2 Results
Table 4 presents the performance of the SQL Agent on 100 curated questions exploring the Venetian cadastral dataset. Performance is evaluated using exact match accuracy, which measures the alignment of generated SQL queries with ground truth, and unigram overlap, which assesses lexical similarity. The ground truth is defined by executing manually annotated SQL queries. We use unigram overlap as a complementary metric because it captures cases where the executed SQL query yields the correct information but differs from the ground truth due to additional contextual details or differences in the order of elements in lists. Over the 100 questions we tried, no SQL runtime errors were observed.
In the zero-shot scenario, the SQL Agent achieves 52% exact match accuracy and 86% unigram overlap. With three-shot prompting, where the model is provided with three example question-query pairs, these metrics improve significantly to 79% and 92%, respectively, illustrating the effectiveness of in-context learning. Despite these strengths, some limitations remain. The system occasionally misinterprets output requirements, such as returning a detailed list instead of an aggregated count for questions like “How many owners receive more than 100 ducati in total rent income?” Similarly, it provides disaggregated results rather than sums for tasks like “What is the total rent income of the top 5 earners?” Complex analytical tasks, such as calculating the “average rent income variance across all locations” or the “share of income from properties labeled as ‘bottega da fabro’,” also pose challenges.
Overall, these results confirm the SQL Agent’s capability to reliably convert natural language queries into executable SQL, enabling effective exploration of historical cadastral data. The high accuracy and minimal errors highlight the utility of text-to-SQL models for idiographic data browsing, while areas requiring further refinement are primarily related to advanced aggregations and formatting nuances.
@@ 5.2 Results
CodeS-7B 0-shot 0.52 0.86
CodeS-7B 3-shot 0.79 0.92
@@
Table 3: Performance of CodeS-7B in the browsing tasks. exact match and unigram overlap scores for 0-shot and 3-shot settings, with zero SQL runtime errors.
6 Prompting cadastre data with text-to-Python
6.1 Related work
The challenge of analyzing historical cadastral data through natural language queries requires bridging several technical domains: data extraction, query interpretation, and automated analysis. Recent advances in AI agents and data-driven discovery provide promising foundations for addressing these challenges. InfiAgent-DABench [17] has established key benchmarks for assessing LLM-based agents in statistical data analysis tasks, particularly highlighting the importance of consistent responses across multiple query iterations. While OpenAgent [18] demonstrates the potential of multi-tool AI agents through combined Python and SQL functionalities, our preliminary experiments revealed that SQL’s rigid structure can be limiting for the dynamic nature of historical data analysis.
Recent work by Kapoor et al. [19] has shifted focus from standardized benchmarks toward optimizing cost-effectiveness and precision in AI agents, informing our approach of limiting debug iterations. The concept of reusable code modules, as demonstrated in CodeChain [20], suggests promising directions for generating adaptable solutions across similar historical queries. Our approach shares commonalities with Majumder et al.’s [21] system in operating across multiple datasets to validate hypotheses, though we extend this to specifically address historical-contemporary comparisons. The capabilities demonstrated by KwaiAgents [22] in performing on-the-fly computations point to future possibilities for incorporating supplementary historical data sources.
Building on these foundations, we present a text-to-Python agent specifically designed for historical cadastral analysis. Our system uniquely combines entity extraction, automated planning, and code generation to bridge historical and contemporary urban data sources. The following subsections detail our method’s implementation and evaluation.
6.2 Method
Questions and data specifications. As discussed in Section 3, the prompting questions fall into four categories: spatial, functional, personal, and temporal. Spatial questions involve locating and organizing datapoints from the Catastici (1740) and Sommarioni (1808) datasets in relation to landmarks derived from the OpenStreetMap-based Landmarks dataset (refer to Figure 5). Temporal questions compare and identify patterns between these datasets. The text-to-Python agent integrates these three datasets for analysis (Figure 5). Critically, this integration enables us to bridge past and present urban landscapes by leveraging semi-persistent landmarks like churches and squares from OpenStreetMap as spatial anchors. While these structures may have evolved physically, their maintained locations and social functions enable meaningful contextualization of historical cadastral data within contemporary geographic frameworks.
Model. The text-to-Python agent functions as a dialogue among three specialized agents: the entity extractor, the planner, and the coder. Each agent is an LLM utilizing specialized prompts (detailed in Supplementary Section A.4). These agents process inputs and produce outputs to guide one another. The overall information flow between agents is managed using Langchain [23] and is illustrated in Figure 5. The following paragraphs detail the implementation and roles of each agent.
Figure 5: The text-to-Python agent. The agent receives a question and consults different datasets to 1) extract the entities being referred to; 2) creates a plan to answer it; and 3) produces and runs a python script to generate an answer.
Entity Extractor. Effective analysis begins with aligning user queries to dataset content. Broad natural language questions often fail to produce precise results due to ambiguous contextual alignment. To address this, a Retrieval-Augmented Generation (RAG) approach [24] integrates dataset values into prompts. For tabular data, our method employs a tailored Entity Extraction phase, which identifies relevant rows and columns to enhance LLM input.
Figure 6 illustrates our system. Upon receiving a query, the system first autonomously selects the appropriate dataset for entity extraction. The "Column Extractor" then aligns the question phrases with corresponding dataset columns, forwarding them to the "Row Extractor". The "Row Extractor" discerns if a phrase denotes a specific term within the column or refers generally to the column. The former undergoes "Entity Search", involving exact matching, fuzzy matching, and semantic search to locate the term within the dataset. Exact matching identifies precise matches (e.g., "avocato" in "Professions"), fuzzy matching finds minor variations (e.g., "avvocato"), while semantic search captures conceptually related terms (e.g., "procuratore"). The three-tier search approach addresses queries that target specific datasets terms. While Semantic Search identifies these, it may also return unintended terms and is resource-intensive. Exact and Fuzzy Matching are thus preferred for their specificity and efficiency.
Code Generation. Inspired by the Plan and Solve [25] approach, the code generation involves two agents: the planner and the coder. the planner creates a detailed solution framework based on the metadata of the data set, the extracted entities, and the query mappings. the coder then translates the plan into executable Python code. an executor runs the code, returning error messages for iterative debugging if needed. after a fixed number of retries, unresolved queries are marked as unanswerable.
Consistency Measures. We evaluate the consistency of the system along the dimension of Execution Consistency (EC), which measures the stability of the system’s responses in three runs with different random seeds.
6.3 Results
Execution Consistency Analysis. Figure 7 illustrates the execution consistency (EC) of answers across three independent runs, analyzed by question category and answer type. We define EC-3 as perfect consistency (identical results across all three runs) and EC-2 as partial consistency (identical results in two out of three runs).
As shown in Figure 7(a), questions about personal information and property functions achieved higher EC scores compared to spatial and comparison queries. Personal queries reached approximately 95% consistency at EC-3, while spatial and comparison queries showed lower consistency at around 85% and 80% respectively. This pattern suggests that straightforward queries about specific entities are more reliably processed than those requiring complex spatial reasoning or comparative analysis.
Figure 7(b) demonstrates that answer format significantly influences consistency levels. Yes/no responses and single numerical answers demonstrated higher consistency (approximately 90% at EC-3) compared to responses requiring entity name extraction (approximately 60% at EC-3). This variation in consistency may be attributed to the more deterministic nature of numerical and binary operations, whereas entity name responses often require multiple processing steps and complex data filtering that can introduce variability across runs.
Our manual verification of EC-3 responses revealed encouraging results: only 12 out of 79 consistent answers contained errors, suggesting that execution consistency serves as a reliable indicator of answer quality. This 15.2% error rate in consistently generated responses indicates that while consistency doesn’t guarantee accuracy, it strongly correlates with correct results.
@@ 6.3 Results
Figure 7: Execution Consistency. grouped by (a) categories and (b) answer types
Qualitative analysis of errors. A key challenge identified pertains to extraction errors resulting from the ambiguous and evolving nature of language. For instance, the terms ’commercial buildings’ or ’shops’ is inherently ambiguous and, within the dataset, it has been erroneously matched to the Venetian term ’magazzeno’ or ’locale’, which typically refers to a garage rather than a commercial space. This lack of specificity further led to imprecise associations, such as the inclusion of the generic term calle when generating a response to the question, “Which square has the largest number of commercial buildings within 100 meters in the dataset in 1740?”
Similarly, diachronic linguistic variations posed significant difficulties, a challenge well-documented in computational historical linguistics [26]. In response to the question, “How many buildings changed from residential to commercial use between 1740 and 1808?”, the system searched for terms such as negozio and ufficio, which are later linguistic evolutions of the historical term bottega. However, these diachronic equivalents are not explicitly mentioned in the cadastre, further complicating the accuracy of the generated response.
Manual verification of all system responses confirmed these patterns of linguistic and contextual inaccuracies, highlighting the need for more sophisticated historical context handling in AI systems. For a detailed analysis of the system’s key data operations and processing methods, see supplementary section B.2. Successful and failing execution traces of the text-to-Python agent are also available in supplementary section B.3.
@@ 7 Discussion
Our experimental results reveal several key insights about the application of LLMs for historical cadastral analysis. First, when comparing the performance of specialized SQL and general-purpose Python agents on structured queries, we find remarkably similar capabilities, with a unigram overlap score of 0.85 for the text-to-Python agent on the questions from Section 5. This suggests that the Python-based approach, while more versatile, does not compromise accuracy on straightforward cadastral queries, offering a potential "one-size-fits-all" solution for various analytical needs.
A fundamental advantage of our text-to-program framework lies in its inherent interpretability and verifiability. By generating executable code rather than direct natural language responses, the system produces solutions that can be traced back to source data, effectively minimizing the risk of hallucinations that often plague LLM applications. Each query result is anchored in specific cadastral records, allowing researchers to verify the exact data points contributing to any conclusion. Furthermore, the generated programs serve as transparent reasoning traces, where analytical assumptions and methodological choices are explicitly encoded in the syntax and logic of the code. Our comparative analysis between GPT-4 [27] and LLama-70B [28], detailed in Supplementary Section B.1, reveals a substantial performance gap between closed-source and open-source models. This difference highlights the current limitations of freely available models for complex historical analysis, though rapid advances in open-source LLMs suggest this gap may narrow in the future. A notable strength of our approach is its city-agnostic nature. While demonstrated on Venetian cadastres, the framework can be readily adapted to any urban area with digitized historical records, requiring only minimal adjustments to accommodate different data structures and local contextual requirements.
However, several limitations must be considered. First, the interpretability of our system presupposes familiarity with programming languages, potentially limiting accessibility for traditional historical researchers. Future work could address this by incorporating an additional layer that translates code into natural language explanations, making the reasoning process as well as the hypotheses made by the system more accessible to non-technical users. Second, our results reveal challenges in handling diachronic language variations (see Section 6.3). This suggests the need for time-aware models with adapted retrieval-augmented generation (RAG) mechanisms capable of processing historical linguistic variations, especially for cases involving Venetian dialect and other historical languages. Third, while our use of semi-permanent urban anchors (such as churches and major squares) effectively grounds historical analysis in modern spatial references, this approach requires careful consideration. One must remain aware of the potential anachronisms and oversimplifications this method might introduce. More broadly, this highlights the importance of understanding the underlying assumptions and limitations of the datasets being analyzed, as these factors significantly influence the interpretation of results generated by the system.
These limitations point to promising directions for future research, including the development of more accessible interpretation tools, enhanced historical language processing capabilities, and more nuanced approaches to temporal-spatial mapping. Despite these challenges, our framework demonstrates the potential of LLM-based approaches to revolutionize historical urban research while maintaining rigorous academic standards through verifiable and interpretable results.

### v2

@@ 4.3 SQL Agent
Architecture. Natural language questions are converted into SQL queries using the open-source text-to-SQL model CodeS-7B li_codes_2024 (23). Each prompt includes a detailed description of the table schema and its columns. We evaluate two prompting settings: in the zero-shot setting, the question is provided with table metadata only; in the three-shot setting, it is provided alongside three example question–query pairs. For each question, the model is run four times with identical inputs, and the final SQL query is selected by majority voting before being executed via SQLite to obtain the corresponding answer. Figure 4 provides an illustration of the SQL Agent. For further clarity, examples of the prompts designed to interact with the CodeS-7B model are provided in Appendix A.2.
4.4 Coding Agents
Questions. As discussed in Section 3, prompting questions fall into four categories: spatial, functional, personal, and temporal. Spatial questions locate and organize records from the Catastici (1740) and Sommarioni (1808) datasets relative to landmarks from the OpenStreetMap-based Landmarks dataset (Figure 5), while temporal questions identify and compare patterns across the two periods. The coding agent integrates these three datasets, using semi-persistent features such as churches and squares as spatial anchors. Although these structures may have changed physically, their stable locations and enduring social functions provide a consistent framework for situating historical cadastral data within contemporary geography.
Figure 5: The coding agent. The agent receives a question and consults different datasets to 1) extract the entities being referred to; 2) creates a plan to answer it; and 3) produces and runs a python script to generate an answer.
Architecture. The coding agent operates as a dialogue among three specialized components: the entity extractor, the planner, and the coder. Each is an LLM prompted for a specific role (see Additional Methods), with outputs from one component guiding the next. The overall information flow is orchestrated using LangChain langchain2022 (6) (Figure 5).
Entity extraction. We use a Retrieval-Augmented Generation (RAG) approach 10.5555/3495724.3496517 (22) to align broad queries with dataset content. The system selects the relevant dataset, maps question terms to columns, and retrieves entries using exact matching (e.g., avocato), fuzzy matching for spelling variants (e.g., avvocato), and semantic matching for related terms (e.g., procuratore). Exact and fuzzy matching offer greater precision, while semantic search increases recall at the cost of specificity (Figure 6).
Code generation. Following the Plan-and-Solve approach wang-etal-2023-plan (30), the planner produces a detailed execution strategy based on dataset metadata, extracted entities, and query mappings. The coder then translates this plan into Python code, which is executed in a controlled environment. Errors trigger iterative debugging up to a fixed retry limit, after which unresolved queries are marked as unanswerable.
Consistency measures. We assess Execution Consistency (EC) as the proportion of identical outputs obtained when repeating the same question three times under different random seeds, capturing the stability of generated code and results. Following prior work kapoor_ai_2024 (19, 17), we use three runs as a cost–accuracy trade-off: it is sufficient to reveal stability patterns while keeping evaluation costs manageable.
5 Results
5.1 SQL Agent Results
We evaluated the SQL Agent on 100 curated browsing questions covering retrieval and relational operations over the Catastici (1740) dataset. Performance was assessed using two complementary metrics: exact match accuracy, measuring the correspondence between generated and ground-truth SQL queries, and unigram overlap, capturing lexical similarity and accounting for cases in which different query formulations yield equivalent outputs. Ground truth queries were manually authored and executed to obtain reference results. No SQL runtime errors were observed in any setting.
In the zero-shot configuration, the SQL Agent achieved 52% exact match accuracy and 86% unigram overlap. Providing three example question–query pairs (three-shot prompting) substantially improved these scores to 79% and 92%, respectively, indicating strong benefits from in-context learning (Table 3).
Error analysis revealed that most failures arose from misinterpreting output specifications. For example, when asked “How many owners receive more than 100 ducati in total rent income?”, the system returned a detailed list rather than an aggregate count. Similarly, in response to “What is the total rent income of the top 5 earners?”, the system provided individual incomes rather than a summed value. More complex analytical tasks, such as computing the “average rent income variance across all locations” or the “share of income from properties labeled as ‘bottega da fabro’,” also proved challenging, reflecting limitations in multi-step aggregation and specialized filtering.
Overall, these results demonstrate that the SQL Agent can reliably translate natural language queries into executable SQL for idiographic browsing of historical cadastral data, with high accuracy and zero execution errors. Remaining weaknesses are primarily associated with advanced aggregation logic and nuanced output formatting. These limitations emerge most clearly when questions require multi-step reasoning, integration of heterogeneous sources, or contextual interpretation—tasks that move beyond structured browsing and into the prompting category addressed by the coding agent.
@@ 5.1 SQL Agent Results
CodeS-7B (0-shot) 0.52 0.86
CodeS-7B (3-shot) 0.79 0.92
@@
Table 3: Performance of CodeS-7B on browsing tasks. Exact match and unigram overlap scores for 0-shot and 3-shot settings, with zero SQL runtime errors.
5.2 Coding Agent Results
Execution consistency. We assessed the coding agent using the Execution Consistency (EC) metric, defined as the proportion of identical answers obtained when executing the same query three times with different random seeds. EC-3 denotes perfect consistency (identical answers in all three runs) and EC-2 denotes partial consistency (identical answers in two out of three runs).
Figure 7 summarizes EC results by question category and by answer type. Personal and functional questions achieved the highest consistency (approximately 95% EC-3), followed by spatial ($\approx$85%) and comparison ($\approx$80%) queries. Analysis by answer type revealed that yes/no and single-numerical responses were more consistent ($\approx$90% EC-3) than entity-name extractions ($\approx$60% EC-3), likely due to the greater determinism of binary and numerical operations compared to multi-step entity retrieval.
Manual inspection of EC-3 responses indicated that 12 of 79 consistently generated answers contained errors, yielding a 15.2% error rate. This finding suggests that while perfect consistency is not a guarantee of correctness, it is a strong indicator of answer reliability.
@@ 5.2 Coding Agent Results
Figure 7: Execution consistency (EC) of the coding agent. (a) EC grouped by question category; (b) EC grouped by answer type.
Model comparison. We compared GPT-4 openai2024gpt4technicalreport (25) and Llama-3 70B grattafiori2024llama3herdmodels (15) on the same prompting tasks (Figure 8). Both models achieved near-perfect execution accuracy (i.e., generated programs ran without error). However, GPT-4 produced markedly higher Execution Consistency and answer correctness. The primary source of this difference was not code syntax, but rather query interpretation during the planning stage. For example, when asked “Which square has the highest density of tenants?”, both models generated valid Pandas code, but GPT-4 correctly mapped “square” to the appropriate geographic entity type and returned campo san bartolomeo, while LLaMA misaligned the entity resolution and returned corte bollani. This illustrates that GPT-4’s advantage stems from emergent reasoning capabilities associated with larger-scale models, enabling more accurate semantic parsing of complex questions, rather than purely from code generation proficiency.
Figure 8: Model comparison on prompting tasks. Performance of GPT-4 and Llama-3 70B in terms of execution consistency and correctness.
5.3 Qualitative Analysis of Data Operations
Table 4 illustrates how the coding agent systematically translates complex natural language questions into executable data analysis operations. Each question is decomposed into key entities and references, mapping linguistic elements (e.g., “rent price,” “square”) to corresponding dataset columns and types. Once entities are identified, the agent determines the appropriate analytical procedure—such as correlation analysis, filtering, or aggregation—tailored to the query. For example, a question about the correlation between rent price and proximity to squares in 1740 is converted into a Pearson correlation calculation, while queries involving temporal comparisons or categorical relationships employ statistical methods such as chi-square tests or comparative metrics.
Two broad error classes emerged from this qualitative inspection. First, semantic ambiguity in entity mapping: modern descriptors such as “commercial building” or “shop” were often mapped to historical terms like magazzeno or locale, which in 18th-century Venetian context usually referred to storage or ancillary spaces. This mismatch inflated counts in spatial aggregation tasks, such as identifying squares with the highest density of commercial buildings. Second, diachronic linguistic variation: terms like negozio or ufficio, introduced in later centuries, were substituted for the historically accurate bottega when analysing functional change over time, leading to systematic undercounting.
From the EC-3 subset of outputs, three consistent trends emerged:
  • Robust mappings in well-defined domains. Queries involving unambiguous spatial anchors (e.g., square, church) or numeric attributes (e.g., rent price) were consistently mapped to correct dataset fields and processed with straightforward statistical or spatial operations, yielding high EC and correctness.
  • Greater variability in functional classifications. Queries involving building function, especially those requiring aggregation over categories (e.g., “multi-function buildings”), were more sensitive to lexical variability and synonym coverage.
  • Temporal integration challenges. Cross-dataset comparisons between 1740 and 1808 often failed to bridge lexical gaps when direct string matches were absent, highlighting the need for historical ontology alignment.
Overall, the coding agent demonstrates strong performance when entity mappings are unambiguous and align with stable schema elements. However, performance degrades when semantic drift, synonym variation, or temporal differences in terminology are present, underscoring the potential value of incorporating structured historical vocabularies into the pipeline.
Question Entity & Reference Code Operation
Is the building rent price correlated with the distance from the nearest square in 1740? “square” $\rightarrow$ type in Landmarks; “rent price” $\rightarrow$ price in Catastici_1740 Pearson correlation between rent price and distance to nearest square
Which neighborhood contains the buildings with the highest rent price on average in 1740? “rent price” $\rightarrow$ price in Catastici_1740 Uses “parish” as “neighborhood”; finds parish with highest average rent price
How many people live within 100 meters from the closest church in 1808? “Church” $\rightarrow$ type in Landmarks; “people” $\rightarrow$ owner name in Sommarioni_1808 Filters buildings within 100 meters of a church; counts unique people
On average, are buildings with multiple functions more expensive than those with a single function in 1740? “single/multiple function” $\rightarrow$ function; “more expensive” $\rightarrow$ price in Catastici_1740 Compares average rent prices for single- vs. multi-function buildings
Which square is surrounded by the most diverse building functions within 100 meters in 1740? “square” $\rightarrow$ type in Landmarks; “diverse function” $\rightarrow$ function in Catastici_1740 Counts building functions within 100 meters; selects square with highest diversity
What is the average distance between workshops and their nearest square in 1740? “workshop” $\rightarrow$ [magazzeno, orto] in function of Catastici_1740 Calculates average distance from workshops to nearest square
Is there a correlation between family name and profession in 1740? “family name”, “profession” $\rightarrow$ profession in Catastici_1740 Creates contingency table; applies chi-square test of independence
In which parish do lawyers own the most buildings in 1740? “parish” $\rightarrow$ parish in Catastici_1740; “lawyers” $\rightarrow$ [avocato, avvocato] in profession Counts buildings by parish; outputs parish with most lawyer-owned buildings
How many medical doctors were there in Venice in 1740? “medical doctor” $\rightarrow$ medico in profession of Catastici_1740 Filters for medico; counts number of doctors
Did the number of multi-function buildings increase from 1740 to 1808? “multi-function” $\rightarrow$ function count in Catastici_1740 and Sommarioni_1808 Compares counts between 1740 and 1808
Which family increased the number of buildings they owned the most from 1740 to 1808? “family” $\rightarrow$ name in Catastici_1740 and Sommarioni_1808 Counts buildings by family in each dataset; outputs family with largest increase
How many new families appeared in Venice in 1808 that were not present in 1740? “new families” $\rightarrow$ name in Catastici_1740 and Sommarioni_1808 Counts families in 1808 not present in 1740
Table 4: Representative examples of EC-3 answers. Each row links question elements to dataset columns and describes the corresponding code operation.
@@ 6 Discussion
Our results provide several insights into the use of LLM-based text-to-program systems for historical cadastral analysis. First, when applied to structured queries, specialized SQL and general-purpose coding agents achieve comparable accuracy: the coding agent reaches a unigram overlap of 0.85 on browsing questions, indicating that its greater versatility does not come at the expense of precision. This suggests that a single Python coding agent pipeline could address both simple and complex queries, offering a unified solution for heterogeneous analytical needs.
A central advantage of the proposed framework is its interpretability. By producing executable code rather than direct natural language answers, the system anchors every result in the source data, reducing the risk of unsupported inferences. The generated programs act as transparent reasoning traces in which analytical assumptions and methodological choices are explicitly encoded. Our comparison of GPT-4 openai2024gpt4technicalreport (25) and Llama-3 70B grattafiori2024llama3herdmodels (15) (Figure 8) shows a substantial performance gap, largely attributable to differences in query interpretation rather than code generation. While this underscores current limitations of open-source LLMs for complex historical analysis, rapid progress in the field suggests the gap may narrow. The framework is also city-agnostic: although demonstrated on Venetian cadastres, it can be adapted to other urban contexts with digitized records, requiring only minor schema and context adjustments.
Several limitations must be considered. First, the interpretability benefits of code outputs assume some programming literacy, which may limit accessibility for historians without technical backgrounds. An intermediate layer translating code into natural language explanations could broaden usability. Second, performance declines when processing diachronic linguistic variation, pointing to the need for time-aware retrieval-augmented generation methods tailored to historical dialects and evolving terminology. Third, the use of semi-permanent urban anchors (e.g., churches, squares) is effective for spatial grounding but may introduce anachronisms or oversimplifications if not critically assessed. More broadly, results must be interpreted in light of the assumptions and biases inherent in the underlying datasets.
These findings highlight promising directions for future work: improving accessibility through natural language code explanations, enhancing historical language processing, and developing more nuanced temporal–spatial integration methods. Despite current challenges, the proposed framework demonstrates that LLM-based agents can support rigorous, verifiable, and scalable historical urban research.
Author Contributions
  • Conceptualization: I. Di Lenardo; F. Kaplan
  • Methodology: T. Karch; J. Saydaliev; I. Di Lenardo
  • Supervision: T. Karch; I. Di Lenardo; F. Kaplan
  • Investigation: T. Karch, J. Saydaliev
  • Data curation: J. Saydaliev
  • Visualization: T. Karch
  • Writing – original draft: T. Karch
  • Writing – review & editing: All authors
All authors approved the final submitted draft.
Competing Interests
The authors declare no competing interests.
Data and Code Availability Statement
An extended version of the 1808 Venetian cadastre is available through a Zenodo repository [^1] https://doi.org/10.5281/zenodo.16761169 di_lenardo_2025_16761169 (11). The exact dataset employed in this study, together with the full set of questions and the accompanying code, are provided in the associated GitHub repository [^2] https://github.com/dhlab-epfl/venice-agents.
Ethical Standards
The research meets all ethical guidelines, including adherence to the legal requirements of the study country.